# MAGELLAN: Metacognitive predictions of learning progress guide autotelic LLM agents in large goal spaces

### v2

@@ Abstract
Machine Learning, ICML
@@ 2.1 Goal selection in autotelic agents
Autotelic agents exploring vast goal spaces face a critical challenge: they must prioritize which goals to pursue to efficiently develop general competence (Colas et al., 2022b). The automatic curriculum learning community has developed various approaches to address this challenge (Portelas et al., 2020b), leveraging different forms of intrinsic motivation: pursuing goals of intermediate difficulty (Held et al., 2017; Racaniere et al., 2020; Castanet et al., 2023), seeking novelty or uncertainty (Warde-Farley et al., 2019; Pong et al., 2019; Pitis et al., 2020), or maximizing learning progress (LP) (Stout & Barto, 2010; Matiisen et al., 2020; Fournier et al., 2018; Portelas et al., 2020a; Colas et al., 2019; Kanitscheider et al., 2021; Kovač et al., 2023; Zhang et al., 2024). Among these, LP-based methods have proven particularly robust — they adapt to the agent’s capabilities without requiring environment knowledge and avoid common pitfalls like getting stuck on goals where progress plateaus or chasing uncontrollable novelty. The key challenge with LP approaches lies in efficiently estimating progress over large goal spaces, which is the focus of our work.
@@ 2.3 Autonomous LLM agents
Recent work has explored using Large Language Models (LLMs) to solve complex tasks in interactive environments. Early approaches focused on direct action prediction using LLMs, either incorporating environmental feedback (Huang et al., 2022; Yao et al., 2022; Hao et al., 2023; Shinn et al., 2023; Wang et al., 2024) or operating without it (Ahn et al., 2022). However, these methods did not update the LLM’s knowledge through environmental interactions. A new direction emerged with GLAM (Carta et al., 2023), followed by (Wen et al., 2024b, a; Zhou et al., 2024) that ground LLMs in interactive environments using online RL. The resulting LLM agents demonstrated remarkable generalization across language tasks; however, they lack the autotelic mechanisms necessary for navigating expansive goal spaces. In this paper, we enhance a SAC-GLAM (Gaven et al., 2024) LLM agent with metacognitive abilities, enabling it to estimate its LP and prioritize goals within large language spaces
@@ 3.3 Classic ALP baselines
We summarize the four methods above and MAGELLAN in Table 1 based on their Efficiency (i.e. computational cost introduced by additional evaluations), Competence Transfer tracking and no Expert Knowledge requirement). We provide in Appendix B the same table for all prior works covered in Section 2.2.
@@
EK-Eval-ALP + $\times$
Eval-ALP $+$ $\checkmark$
EK-Online-ALP $+$ + $\times$
Online-ALP $+$ $\checkmark$
MAGELLAN $+$ $+$ $\checkmark$
@@ 3.4 The Little-Zoo environment as a testbed
Built upon the Playground environment (Colas et al., 2020), Little-Zoo is fully text-based, with observations, goals, and actions expressed in natural language. It features objects that can be combined together and are grouped into the following hidden categories: furniture (which cannot be combined), plants, herbivores, and carnivores. Given the set of all objects and a set of instructions, Little-Zoo’s goal space is the combination of all possible instructions and scene initializations. The feasibility of a goal thus depends on the objects available, making most combinations infeasible and not trivial to detect (see Figure 2 and Appendix A.3). Instructions are hierarchically structured, ranging from simple grasping tasks to more complex sequences involving object interactions (e.g. ”growing” animals). The complete goal space contains approximately 20 million combinations. In our experiments, we subsample goals with the following proportions: 80% of the goals are impossible 16% involve grasping, 3.2% involve growing plants, 0.7% involve herbivores, and 0.1% involve carnivores. These proportions correspond to proportions in the complete goal space (see Appendix A.4).
@@ 4 Experiments
We provide empirical answers to our scientific questions using experiments with 8 different random seeds in the Little-Zoo environment. For our LLM agent, we use SAC-GLAM (Gaven et al., 2024) to finetune Flan-T5 $248$M (Raffel et al., 2020) as per SAC-GLAM’s experiments. We compare MAGELLAN to the classic approaches presented in  3.3. For methods accessing external expert knowledge, we use Little-Zoo’s hidden goal families (grasp any object, grow plant, grow herbivore, grow carnivore) but add only possible goals in these groups. These baselines would totally fail if given all goals from the same hidden family, regardless of their feasibility. We thus provide an additional group containing all impossible goals. In all our experiments, we use success rate (i.e. average outcome over multiple trials for a goal) noted SR as the observed competence.
@@ Acknowledgment
This work was granted access to the HPC resources of IDRIS under the allocation A0171011996 made by GENCI. We acknowledge funding from the European Commission’s Horizon Europe Framework Programme under grant agreement No 101070381 (PILLAR-robots project) and support from the French Defense Innovation Agency (AID). We also thank Nicolas Yax for kindly proofreading the paper and providing valuable feedback.
@@ A.1 Environment mechanics
To interact with objects, multiple actions are accessible. First, the agent can use "Go to {object}" to move to any object. When performing the "Grasp" action, the object the agent is standing on is added to its inventory (up to 2 items). Some objects can interact with each other. To make two objects interact, the agent must release one object in its inventory by using the "Release {object}" action while standing on the other object. Interactions between three objects are also possible: the agent has to hold two objects in its inventory and perform "Release all" while standing on the third object. The ”Release” action is only accessible if an interaction can be made between the object to release and the one the agent is standing on. Therefore, the agent must carefully manage its inventory to avoid filling it with useless objects without being able to release them.
@@ A.3 Example of impossible goals
In our environment, a goal is defined as a combination of objects present in the environment and an instruction, such as ”grasp the rabbit”. Some combinations are incompatible such as:
@@
Another reason is that the object cannot follow the dynamic required in the objective:
  • Growing an object that is neither an animal nor a plant
@@ A.4 Goal repartition
(a) Goal distribution on the categories after generation of the full goal space.
(b) Goal distribution on the categories used in our experiments.
@@ Appendix B Comparison of LP methods
(Oudeyer et al., 2007) $+$ ++ $\checkmark$
(Baranes & Oudeyer, 2009) $+$ ++ $\checkmark$
(Baranes & Oudeyer, 2013) $+$ ++ $\checkmark$
(Moulin-Frier & Oudeyer, 2013) $+$ ++ $\checkmark$
(Moulin-Frier et al., 2014) $+$ ++ $\checkmark$
(Portelas et al., 2020a) $+$ ++ $\checkmark$
(Forestier & Oudeyer, 2016) $+$ ++ $\checkmark$
(Kovač et al., 2023) $+$ ++ $\checkmark$
(Stout & Barto, 2010) $+$ + $\times$
(Lopes & Oudeyer, 2012b) $+$ + $\times$
(Matiisen et al., 2020) $+$ + $\times$
(Fournier et al., 2018) $+$ + $\times$
(Colas et al., 2019) $+$ + $\times$
(Blaes et al., 2019) $+$ + $\times$
(Akakzia et al., 2021) $+$ + $\times$
(Kumar et al., 2024) $+$ + $\times$
(Kanitscheider et al., 2021) $+$ $\checkmark$
(Zhang et al., 2024) $+$ $\checkmark$
@@ Appendix C Implementation details
To facilitate reproduction and future work, we provide our code at https://github.com/LorisGaven/MAGELLAN.
@@ C.4 Compute budget
To optimize GPU VRAM usage during training, we employed 4-bit quantization techniques as described in (Dettmers et al., 2023). We used a vectorized version of Little-Zoo with 32 instances of the environment running (synchronously) in parallel. In order to accelerate training of our LLM agent (both its policy with online RL and MAGELLAN), we leveraged Lamorel [^2] https://github.com/flowersteam/lamorel to deploy 2 instances of the LLM in parallel (distributing both the forward passes to compute actions’ probability or LP and training in a Data Parallelism setting). When using Flan-T5 250M, each LLM instance is distributed (Vertical Model Parallelism) over one Nvidia H100 80GB GPUs requiring thus a total of 2 Nvidia H100 80GB GPUs to run an experiment (1 GPU $\times$ 2 LLM instances). For one seed of one ALP method trained in Section 4.2, performing 500k training episodes requires 80 GPU hours on the Nvidia H100 80GB.
@@ D.1 Ablations on the MAGELLAN architecture
  • A: Different adapters for the policy and MAGELLAN. This is the architecture adopted in the paper. The environment representation for the policy and the goal space representation for the competence estimator are learned separately on different adapters.
  • B: Shared adapters for the policy and MAGELLAN. We test whether a shared useful representation between the policy and competence estimator can emerge through training.
  • C: Shared adapters but only trained with the policy loss. We use a single set of LoRA adapters to train both the policy and the competence estimation module, but apply the latter’s gradient only to the MLP outputting the competence. This ablation tests if the policy loss on its own creates semantical relationships between goals. We argue that training the Q-Value should push towards a space clustering goals.
  • D: Adapter only for the policy, frozen LLM representation for MAGELLAN. In this ablation, we test if MAGELLAN can learn to predict competence using the LLM’s original latent space. Learning to estimate competence over a fixed representation is closer to prior works such as ALP-GMM (Portelas et al., 2020a).
Figure 8: Different architectural choices in MAGELLAN: (a) we learn separate LoRA adapters between the policy and MAGELLAN (used in the paper); (b) we share adapters and update them using both the policy and MAGELLAN’s gradient; (c) we share adapters but they are only updated by the policy’s gradient; (d) MAGELLAN directly uses the latent representation produced by the pretrained LLM.
@@
We compare the learning dynamics of the four architectures in Figure 9. We observe no difference in their ability at learning ”grasp” goals. However, the agents with Architecture A continue to progress on other goal types whereas the ones using Architectures B, C, and D stagnate between 50k and 100k episodes before progressing again. At the end of training (500k episodes), agents using architecture A have learned all goal types. The ones based on Architectures B and C learned up to ”grow herbivore” goal types. The agents that utilize Architecture D succeed plateaued on ”grow plants”.
@@
Architecture A (Figure 10(a)) and architecture B (Figure 10(b)) shared very similar representations with in both cases goals clustered between ”grasp”, ”grow” and impossible goals. Inside the cluster ”grow” the different object types (i.e. plants, herbivores, carnivores) are also clustered. Nonetheless, as seen in Figure 9, Architecture B does not master ”grow carnivore” goals and classified them as impossible. This indicates that obtaining a shared latent space for the policy and MAGELLAN is possible but slows down skill acquisition.

### v3

@@ Abstract
LLM agents, Open-Ended Learning, Learning Progress, Goal-conditionned RL, Automatic Curriculum Learning
@@ 2.1 Goal selection in autotelic agents
Autotelic agents exploring vast goal spaces face a critical challenge: they must prioritize which goals to pursue to efficiently develop general competence (Colas et al., 2022b). The automatic curriculum learning (ACL) community has developed various approaches to address this challenge (Portelas et al., 2020b), leveraging different forms of intrinsic motivation: pursuing goals of intermediate difficulty (Held et al., 2017; Racaniere et al., 2020; Castanet et al., 2023), seeking novelty or uncertainty (Warde-Farley et al., 2019; Pong et al., 2019; Pitis et al., 2020), maximizing regret w.r.t. an optimal policy’s performance  —  mostly used in Unsupervised Environment Design (UED) methods (Dennis et al., 2020; Jiang et al., 2021) —  or maximizing learning progress (LP) (Stout & Barto, 2010; Matiisen et al., 2020; Fournier et al., 2018; Portelas et al., 2020a; Colas et al., 2019; Kanitscheider et al., 2021; Kovač et al., 2023; Zhang et al., 2024). ACL methods also differ in their use of the goal space: while some methods compute metrics over the goal space and sample from it (e.g. (Matiisen et al., 2020; Kanitscheider et al., 2021; Portelas et al., 2020a)), others generate goals (e.g. (Held et al., 2017; Castanet et al., 2023; Dennis et al., 2020)). LP-based methods have proven particularly robust, especially upon a limited training budget in which it is not possible to learn all goals (Lopes & Oudeyer, 2012b). They adapt to the agent’s capabilities without requiring environment knowledge (e.g. which is necessary when using regret) and avoid common pitfalls like getting stuck on goals where progress plateaus or chasing uncontrollable novelty. The key challenge with LP approaches lies in efficiently estimating progress over large goal spaces, which is the focus of our work.
@@ 2.3 Autonomous LLM agents
Recent work has explored using Large Language Models (LLMs) to solve complex tasks in interactive environments. Early approaches focused on direct action prediction using LLMs, either incorporating environmental feedback (Huang et al., 2022; Yao et al., 2022; Hao et al., 2023; Shinn et al., 2023; Wang et al., 2024) or operating without it (Ahn et al., 2022). However, these methods did not update the LLM’s knowledge through environmental interactions. A new direction emerged with GLAM (Carta et al., 2023), followed by (Wen et al., 2024b, a; Zhou et al., 2024) that ground LLMs in interactive environments using online RL. The resulting LLM agents demonstrated remarkable generalization across language tasks; however, they lack the autotelic mechanisms necessary for navigating expansive goal spaces. In this paper, we enhance a SAC-GLAM (Gaven et al., 2024) LLM agent with metacognitive abilities, enabling it to estimate its LP and prioritize goals within large language spaces.
@@ 3.3 Classic ALP baselines
We summarize the four methods above and MAGELLAN in Table 1 based on their Efficiency (i.e. computational cost introduced by additional evaluations), Competence Transfer tracking and no Expert Knowledge requirement. We consider a method’s efficiency as “high” if it does not require any additional evaluation (i.e. it only uses the performance observed on goals sampled), and as “low” otherwise. We evaluate the competence transfer tracking using the following criteria:
  • absence of +: the estimated competence is updated only on sampled goals.
  • +: the estimated competence is updated on a predefined group the sampled goal belongs to.
  • ++: the estimated competence is updated on a dynamically learned group the sampled goal belongs to.
  • $+$: the estimated competence is updated on all goals.
We provide in Appendix B the same table for all prior works covered in Section 2.2.
@@
EK-Eval-ALP low + $\times$
Eval-ALP low $+$ $\checkmark$
EK-Online-ALP high + $\times$
Online-ALP high $\checkmark$
MAGELLAN high $+$ $\checkmark$
@@ 3.4 The Little-Zoo environment as a testbed
Built upon the Playground environment (Colas et al., 2020), Little-Zoo is fully text-based, with observations, goals, and actions expressed in natural language. It features objects that can be combined together and are grouped into the following hidden categories: furniture (which cannot be combined), plants, herbivores, and carnivores. Given the set of all objects and a set of instructions, Little-Zoo’s goal space is the combination of all possible instructions and scene initializations. The feasibility of a goal thus depends on the objects available, making most combinations infeasible and not trivial to detect (see Figure 2 and Appendix A.3). For instance, these are respectively feasible and infeasible goals: “Goal: Grow deer. You see: baby deer, bookshelf, water, tomato seed.”; “Goal: Grow deer. You see: baby deer, bookshelf, baby lion, tomato seed.” (water is missing in the last one). Instructions are hierarchically structured, ranging from simple grasping tasks to more complex sequences involving object interactions (e.g. ”growing” animals). The complete goal space contains approximately 20 million combinations. In our experiments, we subsample goals with the following proportions: 80% of the goals are impossible 16% involve grasping, 3.2% involve growing plants, 0.7% involve herbivores, and 0.1% involve carnivores. These proportions correspond to proportions in the complete goal space (see Appendix A.4).
@@ 4 Experiments
We provide empirical answers to our scientific questions using experiments with 8 different random seeds in the Little-Zoo environment. For our LLM agent, we use SAC-GLAM (Gaven et al., 2024) to finetune Flan-T5 $248$M (Raffel et al., 2020) as per SAC-GLAM’s experiments. We compare MAGELLAN to the classic approaches presented in  3.3. For methods accessing external expert knowledge, we use Little-Zoo’s hidden goal families (grasp any object, grow plant, grow herbivore, grow carnivore) but add only possible goals in these groups. Indeed, these baselines are based on groups predefined in advance by human experts with a strong assumption: the goals within a group share the same learning dynamics and therefore the agent’s competence is the same over all goals in the group. If impossible goals were included, these groups would lose their relevance. Moreover, because of the large number of impossible goals, the average competence within each group will always be very close to 0. There will be no progress niche that the method can use to generate a curriculum, and performance will probably be very close to the random baseline. We thus provide an additional group containing all impossible goals. In all our experiments, we use success rate (i.e. average outcome over multiple trials for a goal) noted SR as the observed competence.
@@ 4.1 How well does MAGELLAN estimate competence (Q1)
To demonstrate that our method generalizes beyond the Little-Zoo environment, we conducted an additional experiment using goals derived from the OpenR1-Math-220k dataset (Face, 2025). In this setup, rather than training a real agent, we simulate the learning of an agent that progressively acquires skills in three categories: Algebra, then Geometry, and finally Number Theory. We compare the competence estimation from MAGELLAN and Online-ALP to the underlying true competence. Figure 4 shows that MAGELLAN accurately tracks the agent’s competence and distinguishes between the different mathematical categories, clearly outperforming Online-ALP. Additional results, including experiments in the BabyAI-Text environment (Chevalier-Boisvert et al., 2019; Carta et al., 2023) and an ablation study on the impact of LLM size within MAGELLAN, are provided in Appendix D.2.
Figure 4: Competence estimation on OpenR1-Math-220k. MAGELLAN (blue) accurately tracks competence across Algebra, Geometry, and Number Theory, closely matching true success probabilities and outperforming Online-ALP (orange).
@@ Acknowledgment
This work was granted access to the HPC resources of IDRIS under the allocation A0171011996 made by GENCI. We acknowledge funding from the European Commission’s Horizon Europe Framework Programme under grant agreement No 101070381 (PILLAR-robots project), the French Defense Innovation Agency, and the ANR AI Chair grant ANR-19-CHIA-0004. We also thank Nicolas Yax for kindly proofreading the paper and providing valuable feedback.
@@ A.1 Environment mechanics
There are $5$ categories of elements the agent can interact with: the objects (i.e. ”bookshelf”), the water, the plants (i.e. tomato) which can bee seed and grown, the herbivores (i.e cow) which can be baby or grown, and the carnivores (i.e. lion) which can be baby or grown.
To interact with any element of the environment, multiple actions are accessible. First, the agent can use "Go to {element}" to move to any element. You can only move to an element (not a random point in the plane) and after a ”Go to” action you are placed on the element. When performing the "Grasp" action, the element the agent is standing on is added to its inventory (up to 2 items). Some elements can interact with each other. To make two elements interact, the agent must release one element in its inventory by using the "Release {element}" action while standing on the other element. Interactions between three elements are also possible: the agent has to hold two elements in its inventory and perform "Release all" while standing on the third element. The ”Release” action is only accessible if an interaction can be made between the element to release and the one the agent is standing on. Therefore, the agent must carefully manage its inventory to avoid filling it with useless elements without being able to release them. At each turn you get a full description of the environment with "You see {element_1}, {element_2}, {element_3}, {element_4}".
@@ A.3 Example of impossible goals
In our environment, a goal is defined as a combination of elements present in the environment and an instruction, such as ”grasp the rabbit”. Some combinations are incompatible such as:
@@
Another reason is that the element cannot follow the dynamic required in the objective:
  • Growing an element that is neither an animal nor a plant
@@ A.4 Goal repartition
(a) Goal distribution per categories on the full goal space.
(b) Goal distribution per categories in our experiments.
Figure 8: We first generate the full goal space whom the distribution is given Figure 8(a) then for computational reasons we sample a smaller goal space with the following distribution Figure 8(b).
@@ Appendix B Comparison of LP methods
We consider a method’s efficiency as “high” if it does not require any additional evaluation (i.e. it only uses the performance observed on goals sampled), and as “low” otherwise. We evaluate the competence transfer tracking using the following criteria:
  • absence of +: the estimated competence is updated only on sampled goals.
  • +: the estimated competence is updated on a predefined goal subset the sampled goal belongs to.
  • ++: the estimated competence is updated on a dynamically learned goal subset the sampled goal belongs to.
  • $+$: the estimated competence is updated on all goals.
@@
(Oudeyer et al., 2007) high ++ $\checkmark$
(Baranes & Oudeyer, 2009) high ++ $\checkmark$
(Baranes & Oudeyer, 2013) high ++ $\checkmark$
(Moulin-Frier & Oudeyer, 2013) high ++ $\checkmark$
(Moulin-Frier et al., 2014) high ++ $\checkmark$
(Portelas et al., 2020a) high ++ $\checkmark$
(Forestier & Oudeyer, 2016) high ++ $\checkmark$
(Kovač et al., 2023) high ++ $\checkmark$
(Stout & Barto, 2010) high + $\times$
(Lopes & Oudeyer, 2012b) high + $\times$
(Matiisen et al., 2020) high + $\times$
(Fournier et al., 2018) high + $\times$
(Colas et al., 2019) high + $\times$
(Blaes et al., 2019) high + $\times$
(Akakzia et al., 2021) high + $\times$
(Kumar et al., 2024) high + $\times$
(Kanitscheider et al., 2021) low $+$ $\checkmark$
(Zhang et al., 2024) low $+$ $\checkmark$
@@ Appendix C Implementation details
To facilitate reproduction and future work, we provide our code at https://github.com/flowersteam/MAGELLAN.
@@ C.4 Compute budget
To optimize GPU VRAM usage during training, we employ 4-bit quantization techniques as described in (Dettmers et al., 2023). We use a vectorized version of Little-Zoo with 32 instances of the environment running (synchronously) in parallel. In order to accelerate training of our LLM agent (both its policy with online RL and MAGELLAN), we leverage Lamorel [^2] https://github.com/flowersteam/lamorel to deploy 2 instances of the LLM in parallel. Thus, we distribute both the forward passes to compute actions’ probability or LP and training in a Data Parallelism setting. When using Flan-T5 250M, each LLM instance is distributed (Vertical Model Parallelism) over one Nvidia H100 80GB GPUs requiring thus a total of 2 Nvidia H100 80GB GPUs to run an experiment (1 GPU $\times$ 2 LLM instances). For one seed of one ALP method trained in Section 4.2, performing 500k training episodes requires 80 GPU hours on the Nvidia H100 80GB.
@@ D.1 Ablations on the MAGELLAN architecture
  • A: Different adapters for the policy and MAGELLAN (Figure 9(a)). This is the architecture adopted in the paper. The environment representation for the policy and the goal space representation for the competence estimator are learned separately on different adapters.
  • B: Shared adapters for the policy and MAGELLAN(Figure 9(b)). We test whether a shared useful representation between the policy and competence estimator can emerge through training.
  • C: Shared adapters but only trained with the policy loss(Figure 9(c)). We use a single set of LoRA adapters to train both the policy and the competence estimation module, but apply the latter’s gradient only to the MLP outputting the competence. This ablation tests if the policy loss on its own creates semantical relationships between goals. We argue that training of the policy should push towards a clustering of the goal space.
  • D: Adapter only for the policy, frozen LLM representation for MAGELLAN(Figure 9(d)). In this ablation, we test if MAGELLAN can learn to predict competence using the LLM’s original latent space. Learning to estimate competence over a fixed representation is closer to prior works such as ALP-GMM (Portelas et al., 2020a).
Figure 9: Different architectural choices in MAGELLAN: (a) we learn separate LoRA adapters between the policy and MAGELLAN (used in the paper); (b) we share adapters and update them using both the policy and MAGELLAN gradient; (c) we share adapters but they are only updated by the policy gradient; (d) MAGELLAN directly uses the latent representation produced by the pretrained LLM.
@@
We compare the learning dynamics of the four architectures in Figure 10. We observe no difference in their ability at learning ”grasp” goals. However, the agents with Architecture A continue to progress on other goal types whereas the ones using Architectures B, C, and D stagnate between 50k and 100k episodes before progressing again. At the end of training (500k episodes), agents using architecture A have learned all goal types. The ones based on Architectures B and C learned up to ”grow herbivore” goal types. The agents that utilize Architecture D succeed plateaued on ”grow plant”.
@@
Architecture A (Figure 11(a)) and architecture B (Figure 11(b)) shared very similar representations with in both cases goals clustered between ”grasp”, ”grow” and impossible goals. Inside the cluster ”grow” the different element types (i.e. plants, herbivores, carnivores) are also clustered. Nonetheless, as seen in Figure 10, Architecture B does not master ”grow carnivore” goals and classified them as impossible. This indicates that obtaining a shared latent space for the policy and MAGELLAN is possible but slows down skill acquisition.
@@ D.2.1 Per-goal competence estimation
D.2.2 Competence estimation on the BabyAI-text goal-space
We replicated the experiment from Section 4.1, simulating a learning agent and estimating its competence online using MAGELLAN and Online-ALP. In this setting, the goal space is drawn from the BabyAI-Text environment, consisting of five goal categories of increasing difficulty: Go to ”object”, Pick up ”object”, Open ”door”, Put ”object A” next to ”object B”, and Pick up ”object A” then go to ”object B”. As shown in Figure 15, MAGELLAN outperforms Online-ALP, successfully tracking the agent’s competence progression across the different task categories.
Figure 15: Competence estimation on BabyAI-Text. MAGELLAN accurately tracks competence across five goal types of increasing difficulty and consistently outperforms Online-ALP.
D.2.3 Impact of the LLM used in MAGELLAN
In this section, we compare different LLMs (Flan-T5-small, Flan-T5-base, and Qwen2.5-0.5B (Qwen et al., 2025)) to evaluate their impact on the competence estimation performance of MAGELLAN. The experiments simulate agent learning using two goal spaces: OpenR1-Math-220k and BabyAI-Text (Figures 16 and 17). Results show that all tested LLMs achieve similar performance, suggesting that MAGELLAN’s competence estimation remains robust across a range of lightweight language models.
Figure 16: Effect of LLM choice on MAGELLAN’s competence estimation for OpenR1-Math-220k. All models yield similar performance.
Figure 17: Effect of LLM choice on MAGELLAN’s competence estimation for BabyAI-Text. Performance remains consistent across models.